# Unclonable Encryption with Continuous Variables

### v1

@@ 4.4 Monogamy of entanglement game
Let $w^{\rm UE}(N,t)$ denote the winning probability for Game 4.3. Let $w^{\rm full}(N-2t)$ be the winning probability for Game 4.4 with $\delta=\alpha$, $\varepsilon=\alpha$, $M=N-2t$. It holds that
@@
Next, in Game 4.3 it is required to get at least $N-t$ bits of $c$ correct. Let $e\in\{0,1\}^{N}$ denote an error pattern. Let $w^{\rm UE}_{e}$ be the probability of winning with precisely error pattern $e$. The bit errors can be arbitrarily distributed over the $q$ part and the $p$ part, which does not nicely fit the symmetric structure of the monogamy game. In order to obtain the symmetric structure we loosen our requirements a little, and allow a surplus of bit errors in one block (if any) to be balanced by additionally allowed bit errors in the other block. In the worst case, all bit errors are located in one block; this leads to an allowed $2t$ errors. We write
@@ 4.5 Main theorem
Note that asymptotically $N$ gets close to $\frac{n}{1-h(\beta)}$, where $h$ is the binary entropy function and $\beta=\frac{1}{2}{\rm Erfc}(\alpha\sqrt{\cosh r})$ is the bit error rate specified in the proof of Theorem 3.1. We see from the result (15) that $\alpha$ needs to be smaller than approximately $\frac{1}{2}$, otherwise the attackers’ winning probability is not exponentially small in $n$. This is unsurprising, for the following reason. One obvious cloning attack on our scheme is to perform a heterodyne measurement on each cipherstate mode individually. This corresponds to mixing the squeezed state with the vacuum state using a 50/50 beam splitter. The effect is a deterioration of the signal-to-noise ratio: On the one hand, the signal power $\alpha^{2}$ is reduced to $\alpha^{2}/2$. On the other hand, the noise power goes from $\frac{1}{2\cosh r}$ to $\frac{1}{2}(\frac{1}{2}+\frac{1}{2\cosh r})$ due to the averaging with the vacuum’s shot noise, which has power $\frac{1}{2}$. The resulting signal-to-noise ratio in the heterodyne measurement is $2\alpha^{2}/(1+\frac{1}{\cosh r})$ $\approx 2\alpha^{2}$. Hence at $\alpha={\cal O}(1)$ there is significant leakage about the codeword bits $c_{i}$, putting the unclonability at risk.
It is interesting to note that the precise value of the squeezing parameter $r$ has little influence on the unclonability property. The $r$ has to be set such that the bit error rate $\beta$ for the legitimate recipient is manageable, so that decryption succeeds with very high probability. I.e. the signal-to-noise ratio needs to be large enough. Other than that, there are no constraints on $r$.
Fig. 2 plots the winning probability of our UE game as a function of message size, including also the UE game for conjugate coding into qubit states [broadbent_et_al2020]. The conjugate coding scheme first one-time pads a message $m$ and then bitwise encodes the ciphertext into the standard basis or the Hadamard basis. Note that Fig. 2 is just an illustration for a single choice of parameters, and is not representative. Furthermore, a comparison of DV and CV is not as straightforward as directly comparing the winning probability in the cloning game.
Figure 2: Winning probabilities in the unclonability game for various schemes. The ideal curve is the simple guessing probability $\left(\frac{1}{2}\right)^{n}$. The wining probability of the conjugate coding scheme is at most $\left(\frac{1}{2}+\frac{1}{2\sqrt{2}}\right)^{n}$. For our CV scheme we plotted (15) with parameters $r=12,t=1$ and $\alpha=0.01$.
@@ 5 Discussion
We have introduced the first Unclonable Encryption scheme with Continuous-Variable states, and have given a proof of unclonability in the game-based framework of [broadbent_et_al2020]. This brings an interesting cryptographic primitive into the CV domain, making it potentially cheaper to implement.
We see several avenues for improvement. It is possible to get some extra tightness in the inequalities. The $\tau(\lambda)$ can be slightly improved by performing error correction on the $q$-block and $p$-block separately; this would change the $N-2t$ in (13) to $N-t$.
@@
As a topic for future work we mention tolerance to channel noise. Channel noise will have a rather severe impact, since attenuation reduces the squeezing. In DV a similar effect arises: photon loss necessitates the use of error-correcting codes that can deal with erasures, which has a crippling effect on the efficiency of UE.
@@ Appendix A The Entangled State
Now we switch to the displaced EPR state, where both modes are translated over ±$\pm\alpha$ in the $\phi=0$ direction,
@@
$\left|{\rm EPR}_{r,0}(\pm\alpha)\right\rangle_{\rm restr}$ $\propto$ ±$\int_{\pm\alpha-\alpha}^{\pm\alpha+\alpha}\!{\rm d}x_{1}\;\left|x_{1}\right\rangle\otimes\int_{-\infty}^{\infty}\!{\rm d}x_{2}\;\left|x_{2}\right\rangle\langle x_{1},x_{2}|\text{EPR}_{r,0}(\pm\alpha)\rangle$ (20)
$\propto$ ±$\!\!\int_{\pm\alpha-\alpha}^{\pm\alpha+\alpha}\!\!\!\!{\rm d}x_{1}\;\left|x_{1}\right\rangle e^{-\frac{(x_{1}-\pm\alpha)^{2}}{2\cosh r}}\otimes\int_{-\infty}^{\infty}\!\!\!\!{\rm d}x_{2}\;\left|x_{2}\right\rangle e^{-\frac{\cosh r}{2}[x_{2}-\pm\alpha-(x_{1}-\pm\alpha)\tanh r]^{2}}\quad\quad$ (21)

### v2

@@ 4.4 Monogamy of entanglement game
Let $w^{\rm UE}(N,t)$ denote the winning probability for Game 4.3. Let $w^{\rm full}(N-2t)$ be the winning probability for Game 4.4 with $\delta=\alpha$, $\varepsilon=\alpha$, M=N2tM=N-2t. It holds that
@@
Next, in Game 4.3 it is required to get at least NtN-t bits of $c$ correct. Let $e\in\{0,1\}^{N}$ denote an error pattern. Let $w^{\rm UE}_{e}$ be the probability of winning with precisely error pattern $e$. The bit errors can be arbitrarily distributed over the $q$ part and the $p$ part, which does not nicely fit the symmetric structure of the monogamy game. In order to obtain the symmetric structure we loosen our requirements a little, and allow a surplus of bit errors in one block (if any) to be balanced by additionally allowed bit errors in the other block. In the worst case, all bit errors are located in one block; this leads to an allowed $2t$ errors. We write
@@ 4.5 Main theorem
We see from the result (16) that $\alpha$ needs to be smaller than approximately $\frac{1}{2}$, otherwise the attackers’ winning probability is not exponentially small in $n$. This is unsurprising, for the following reason. One obvious cloning attack on our scheme is to perform a heterodyne measurement on each cipherstate mode individually. This corresponds to mixing the squeezed state with the vacuum state using a 50/50 beam splitter. The effect is a deterioration of the signal-to-noise ratio: On the one hand, the signal power $\alpha^{2}$ is reduced to $\alpha^{2}/2$. On the other hand, the noise power goes from $\frac{1}{2\cosh r}$ to $\frac{1}{2}(\frac{1}{2}+\frac{1}{2\cosh r})$ due to the averaging with the vacuum’s shot noise, which has power $\frac{1}{2}$. The resulting signal-to-noise ratio in the heterodyne measurement is $2\alpha^{2}/(1+\frac{1}{\cosh r})$ $\approx 2\alpha^{2}$. Hence at $\alpha={\cal O}(1)$ there is significant leakage about the codeword bits $c_{i}$, putting the unclonability at risk.
We consider the asymptotic case $n\to\infty$. In this limit the code rate approaches capacity and we write $n\approx N[1-h(\beta)]$. From the security definition we have the requirement that $2^{-n+\tau}$ should be asymptotically small. Substituting $n\to N[1-h(\beta)]$ in (16), the requirement $n+\tau<0$ can be rewritten as
$h(\beta)-(\frac{1}{2}-\beta)[1-\log(1+2\alpha)]<0.$ (17)
As $\beta$ is a function of $\alpha$ and $r$ (12), eq.(17) specifies a region in the $(\alpha,r)$-plane where it is possible to achieve security asymptotically.
Figure 2: We plot (17) as a function of the squeezing $r$ and displacement $\alpha$. As illustrated, asymptotically the scheme is secure only for $r>3.5$ for some values of $\alpha$.
5 Practical considerations
Since our main motivation for constructing CV-UE is to have a practically implementable scheme, we examine how the protocol behaves under realistic parameter choices and noise.
Generating highly squeezed light is experimentally challenging. We therefore restrict $r$ to values that are reasonably close to the current state of the art. A $15$ dB squeezed vacuum has been demonstrated [Vahlbruch2016], and $10$ dB is common.11110 dB corresponds to $e^{r}=10$, $r\approx 2.30$; 15 dB corresponds to $e^{r}=10^{1.5}\approx 31.6$, $r\approx 3.45$.
The bit error rate (BER) $\beta$ is given by (12). We aim for a small BER, which in turn yields a small decryption–failure probability $\varepsilon_{\rm DF}$ (11). Fig. 3(a) displays $\beta$ as a function of $r$, at $\alpha=0.4$.
As an example, consider $\alpha=0.40$, $r=3.4$, yielding $\beta=0.014$. We take a ciphertext of size $N=1000$, and we set $t=35$ so that $t>N\beta$. Then (11) gives $\varepsilon_{\rm DF}=6.9\cdot 10^{-6}$. Note that the actual message size in this case will be close to $N(1-h(\beta))\approx 892$. We will continue to use this relationship between $N$ and $n$ in the remainder of this section.
(a) $\beta_{\mathrm{noisy}}$ as a function of squeezing parameter $r$, at displacement $\alpha=0.4$ plotted for various transmittance $T$. Excess noise $\xi=0.001$ for all curves.
(b) $\beta_{\mathrm{noisy}}$ as a function of channel transmittance $T$ and excess noise $\xi$. Parameters: $r=3.6$, $\alpha=0.4$.
Figure 3: Bit error rate (19) for different parameters (channel parameters in SNU).
5.1 Robustness under noise
We model the optical channel by its transmittance $T\in[0,1]$ and (input-referred) excess-noise power $\xi$ in SNU [leverrier:tel-00451021]. In this model, the receiver’s effective quadrature variance and displacement transform as
$\frac{1}{\cosh r}\longrightarrow\frac{T}{\cosh r}+(1-T)+T\xi,\qquad\alpha\longrightarrow T\alpha,$ (18)
which follows from the standard covariance–matrix transformation of a thermal–loss (attenuator) Gaussian channel and linear detection [garcia-patron-sanchez2007thesis]. Detector inefficiency and electronics noise can be absorbed into the overall transmittance/noise budget in this framework [YMKSK2023]. Under the map (18), the BER becomes
$\beta_{\mathrm{noisy}}=\tfrac{1}{2}\mathrm{Erfc}\left(\frac{T\alpha}{\sqrt{\dfrac{T}{\cosh r}+(1-T)+T\xi}}\right).$ (19)
This expression is plotted in Fig. 3(b) as a function of $(T,\xi)$ for fixed $(r,\alpha)$.
As an example, consider 5 km of standard single-mode optical telecom fibre at $1550$ nm wavelength. The loss rate is $0.22$ dB/km; at $5$ km this corresponds to $T\approx 0.8$. At $15$ dB squeezing ($r\approx 3.5$), Fig 3(a) then gives a BER around 2020%, which can be handled with a low-rate code, e.g. an LDPC code, or a higher-rate code concatenated with a repetition code.
Note that we assume that the adversaries are not using the noisy channel; they have $T=1$ and $\xi=0$. The channel properties hence need not enter the security analysis.
@@ 6 Discussion
We have introduced the first Unclonable Encryption scheme with Continuous-Variable states, and have given a proof of unclonability in the game-based framework of [broadbent_et_al2020]. To the best of our knowledge, this is also the first analysis of an unclonable encryption protocol in the presence of realistic channel noise, marking a significant step toward practical quantum cryptography.
Making a meaningful comparison with prior work is difficult, since DV–CV performance cannot be fairly assessed by directly comparing the winning probabilities in the cloning game. The most informative benchmark is probably the DV analogue of our scheme, which can be realized via Wiesner’s conjugate coding [Wiesner1983]. Fig. 4 plots the winning probability of our UE game as a function of message size, including also the UE game for conjugate coding into qubit states [broadbent_et_al2020]. The conjugate coding scheme first one-time pads a message $m$ and then bitwise encodes the ciphertext into the standard basis or the Hadamard basis. For experimentally feasible squeezing of $r=3.6(\approx 15.6\;\mathrm{dB})$, we observe exponentially decreasing probability of the adversaries winning only for message longer message sizes in the order of $10^{2}$ or greater. We emphasize that Fig. 4 is just an illustration for a single choice of parameters, and is not intended as a comprehensive performance comparison.
Figure 4: Winning probabilities in the unclonability game for various schemes. The ideal curve is the simple guessing probability $\left(\frac{1}{2}\right)^{n}$. The wining probability of the conjugate coding scheme is at most $\left(\frac{1}{2}+\frac{1}{2\sqrt{2}}\right)^{n}$. For our CV scheme we plotted $2^{-n+\tau(\lambda)}$ (16) with parameters $r=3.6,\alpha=0.4$ and allow bit flips $t=3.5\%$ of ciphertext size.
We see several avenues for improvement. It is possible to get some extra tightness in the inequalities. The $\tau(\lambda)$ can be slightly improved by performing error correction on the $q$-block and $p$-block separately; this would change the N2tN-2t in (14) to NtN-t.
@@
As discussed above, our results establish the security notion of unclonability, rather than the stronger guarantee of unclonable indistinguishability. Arguments based on monogamy of entanglement do not appear to yield guarantees beyond search security; establishing unclonable indistinguishability for CV protocols will likely require new proof techniques. We believe this is attainable and leave a full treatment to future work.
@@ Appendix A The Entangled State
Now we switch to the displaced EPR state, where both modes are translated over $\pm\alpha$ in the $\phi=0$ direction,
@@
$\left|{\rm EPR}_{r,0}(\pm\alpha)\right\rangle_{\rm restr}$ $\propto$ $\int_{\pm\alpha-\alpha}^{\pm\alpha+\alpha}\!{\rm d}x_{1}\;\left|x_{1}\right\rangle\otimes\int_{-\infty}^{\infty}\!{\rm d}x_{2}\;\left|x_{2}\right\rangle\langle x_{1},x_{2}|\text{EPR}_{r,0}(\pm\alpha)\rangle$ (24)
$\propto$ $\!\!\int_{\pm\alpha-\alpha}^{\pm\alpha+\alpha}\!\!\!\!{\rm d}x_{1}\;\left|x_{1}\right\rangle e^{-\frac{(x_{1}-\pm\alpha)^{2}}{2\cosh r}}\otimes\int_{-\infty}^{\infty}\!\!\!\!{\rm d}x_{2}\;\left|x_{2}\right\rangle e^{-\frac{\cosh r}{2}[x_{2}-\pm\alpha-(x_{1}-\pm\alpha)\tanh r]^{2}}\quad\quad$ (25)